Place parentheses to make the equation 6+12-5*3=27 true.

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

6+(12-5)*3=27
6+(7*3)=27
6+21=27

Answer by solver91311(24713)



Hint: 3 times 7 is 21

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism